OpenStudy (anonymous):

what is the x and y intercept of 7x -7y =35

OpenStudy (stefrheart):

In order to find the x intercept replace the y with 0 and to find the y intercept replace the x with 0

OpenStudy (anonymous):

I knew how to do that I just wasn't sure about the signs (negative or positive)

OpenStudy (anonymous):

Find the x and y intercept for the following equation 7x -7y =35 ( I know to plug in x, im just not sure of the signs)

OpenStudy (triciaal):

x intercept when y = 0 so 7 x- 7(0) = 35 x = 35/7 = 5 y intercept when x = 0 so 7(0) - 7 y = 35 y = 35/-7 = -5